WACC Kose, Inc., has a target debt-equity ratio of .38. Its WACC is 10.1 percent and...

WACC

Kose, Inc., has a target debt-equity ratio of .38. Its WACC is 10.1 percent and the tax rate is 25 percent.

a. If the company’s cost of equity is 12 percent, what is its pretax cost of debt?

b. If instead you know that the after tax cost of debt is 6.4 percent, what is the cost of equity?
Answer #1

WACC = kex We + KdCl-t) Wa • Debtl Equity Ratio = 0,38 10.17. - 127.X 1_+ kd (1-0-25)¥ 0.38 1.38 1.38 → 10.1 = 8.69565 + 0.20(6) WACC = kex wet kdCl-t) Wd 10.1% = ke x 1 2 + 6. W x 0.38 1.38 = 10.1 = 0.724637 ke +1-7623188 Ke = 10.1 -11.7623 1884 0.7